OpenStudy (anonymous):

find the critical numbers: t^4+t^3+t^2+1

OpenStudy (agreene):

\[\frac{d}{dt} (t^4+t^3+t^2+1)= t (2+3 t+4 t^2)\] \[t (2+3 t+4 t^2)=0\] \[t= 0; \space \space t=\pm\frac18i(\sqrt{23}+3i)\]